OpenStudy (anonymous):

1.) Calculating the future value of property: Ben Collins plans to buy a house for $220,000. If that real estate is expected to increase in value by 3 percent each year, what will its approximate value be seven years from now?

OpenStudy (anonymous):

After seven years of interest, Ben will pay about $263,479.25. ??

OpenStudy (anonymous):

220,000×.03=6,600 220,000 + 6,600 = 226,600 226,000×.03=6,798 226,600+6,798=227,2798 227,2798×.03=6,818.94 227,2798 + 6,818.94 = 234,097.94 234,097.94×.03=7,022.91 234,097.94+7,022.91=241,120.85 241,120.85×.03=7,233.62 241,120.85+7,233.62=248,354.47 248,354.47×.03=7,450.63 248,354.47+7,450.63=255,805.10 255,805.10×.03=7,674.15 255,805.10 + 7,674.15 =263,479.25 Final answer: $263, 479.25

OpenStudy (anonymous):

\[A=P(1+i)^{n}\] \[A=220000(1.03)^{7}\] A=270572.2503934714 A = $270,572.25